# Compensation behaviour in trilayered anisotropic 6-state clock model

### v1

@@ Abstract
The equilibrium behaviours of a trilayered 6-state clock model has been investigated by Monte Carlo simulation. The intralayer interaction is considered ferromagnetic, whereas the interlayer interaction is antiferromagnetic. Open boundary conditions are applied in all directions. The thermodynamic behaviours of sublattice magnetisation, total magnetisation, magnetic susceptibility and the specific heat are studied as functions of the temperature. The interesting compensation phenomenon, i.e., the vanishing of the total magnetisation even with nonvanishing sublattice magnetisation, has been observed. The compensation temperature and the critical temperature are found to depend on the strength of single-site anisotropy. The comprehensive phase diagram is shown in the temperature-anisotropy plane. The finite size analysis has been done through the temperature dependence of fourth order Binder cumulant, and the scaling exponents are estimated.
@@ 1 Introduction
The ferromagnetic models showing the unusual ordered phase even in absence of conventional long-range ferromagnetic order is a remarkable discovery[1, 2]. The continuous symmetric (SO(2)) planar ferromagnetic model shows this behaviour in two dimensions with vortex-antivortex special kind of ordered state[3, 4, 5]. The unconventional metastable behaviour and phase transition in this kind of model has drawn the attention of the researchers in last four decades. The celebrated BKT transition is now under the spotlight of modern research of condensed matter physics and statistical mechanics.
@@
Is this kind of unusual phase the property of continuous symmetry only? This question has triggered to investigate the discrete symetric q-state clock model, where the continuous symetric XY model can be restored in the limit $q\to\infty$. The properties of q-state clock model has been investigated[6] for q=4, 5 and 6 in two dimensions by Monte Carlo renormalization group analysis to study the phase transitions. The Kosterlitz-Thouless type intermediate phase has been found[7] in two dimensional 6-state clock model by Monte Carlo simulations. The KT type phase transition has been found to occur in the range between ${{k_{B}T}\over{J}}=0.68\pm 0.02$ and ${{k_{B}T}\over{J}}=0.92\pm 0.01$. However, it is confirmed later that there exists a Kosterlitz-Thouless (KT) like phase in the ferromagnetic 6-clock model on the square lattice through studies of the interfacial free energy estimated by Monte Carlo simulations and authors estimated[8] that lower and upper transition temperatures are 0.75 and 0.90 respectively. Interestingly, in a recent study[9], the machine-learning detection of BKT transition was found in q-state clock model with the neural network based estimates of the two critical points as ${{k_{B}T}\over{J}}=0.410$ and ${{k_{B}T}\over{J}}=0.921$.
@@
The classical 2D clock model is known to have a critical phase with Berezinskii–Kosterlitz–Thouless (BKT) transitions. These transitions have logarithmic corrections which make numerical analysis difficult. Hence, the classical 2D six-state clock model has been studied[10] by mapping it to the quantum self-dual 1D six-state clock model and confirmed that the self-dual point has a precise numerical agreement with the analytical result. It has been argued also that the degeneracy of the excited states at the self-dual point from the effective field theoretical point of view.
@@
The thermodynamic behaviours of the clock model has been studied in digfferent kinds of lattices to investigate the effects of geometrical connections and interactions. The BKT transition has been observed[11] in 6-state clock model on the rewired squared lattice studied by Monte Carlo simulation with Wang-Landau algorithm and estimated the lower and upper critical temperatures. For pure case, i.e., the lattice with average number of bonds equal to 2, ${k_{B}T1}\over{J}$} and ${k_{B}T2}\over{J}$}, are roughly estimated around 0.55 and 1.15 respectively (these results may be compared to those reported in the references [7] [8] and [9]). The BKT transition has been studied[12] recently, in 6-state clock model on different kinds (triangula, kagome, honyecomb and diced) of two dimensional lattices. The two dimensional q-state clock model on random lattices has been investigated[13] via numerical simulation. Interestingly, for q=4 a single phase transition has been reported. Whereas, for q=5 and q=6 multiple (double) phase transitions are observed from the two peaks of the specific heat. The lower and upper critical temperatures are found to depend significantly on the structure of the lattice. The anisotropic (with anisotropy parameter $\alpha$) 6-state clock model has been investigated[14] by transfer matrix renormalization group method to explore the phase transition in two dimensional square lattice. Three distinct phase are identified and a comprehensive phase diagram is drawn. They observed a single discontinuous phase transition and two BKT phase transitions meeting in the tricritical point. The tricritical point is estimated at critical anisotropy $\alpha_{c}=0.21405$. The antiferromagnetic 5-state clock-model has been studied on the triangular lattice recently[15] with second nearest neighbour interaction and the phase diagram is drawn.
@@
The thermodynamic behaviours of clock model in layered structured lattices are not yet reported in the well developed literature of this specific model. However, the behaviours of layered Ising and Blume-Capel models are well studied. Let us briefly mention a few selected studies on layered Ising and Blume-Capel models. The main interesting phenomenon in the thermodynamic behaviours of layered Ising and Blume-Capels models is the compensation transition. The compensation temperatures and magnetic susceptibility of a mixed ferro-ferrimagnetic ternary alloy has been studied[16] by Monte Carlo simulations and observed outstanding behaviour of magnetic susceptibility near the compensation and critical temperature. The compensation temperature has been found[17] to depend of the anisotropy in a mixed spin (spin-${1}\over{2}$} and spin-$1$) Heisenberg ferrimagnet on a simple cubic lattice, studied by Oguchi approximation. The anisotopy dependent spin compensation temperature has also been noticed[18] in mixed spin (spin-${3}\over{2}$} and spin-${5}\over{2}$}) ferrimagnetic Ising model studied by mean field approximation. In this case, the possibility of multiple compensation temperature was indicated. The impurity- concentration dependent compensation[19] and the anisotropy dependent compensation[20] was also noticed in the trilayered Ising and Blume-Capel models, respectively. The meanfield and effective field theory has been employed[21] to study the interaction dependent compensation in spin-${1}\over{2}$} Ising trilayered ferrimagnetic system. The phase diagram (compensation and no-compensation) aws drawn in parameter space. The antiferromagnetic systems showed the compensation behaviours in Ising mixed spin (spin-${3}\over{2}$} and spin-${5}\over{2}$}) core/shell nanowire[22] and in mixed spin (spin-${7}\over{2}$} and spin-1{1}) antiferromagnetic ovalene nanowire[23]. The compensation wad found[24] Ising mixed spin (spin-${3}\over{2}$} and spin-${1}\over{2}$}) graphene bilayered ferrimagnet. The effects of exchange interactions, crystal field and external magnetic field on the compensation temperature was studied[24] by Monte Carlo simulation. The evidence of engineering the compensation in ferrimagnetic iron garnet by tuning the ratio of in-plane and out of plane magnetisations can be found in [25]. An interesting spin compensation and first order reentrant behaviour has been observed[26] in mixed spin (spin-1 and spin-2) Ising ferrimagnetic system studied by Monte Carlo simulation.
@@
Can the discrete symmetric planar models (6-state clock model) show the compensation behaviour? Being motivated by this question, we have thoroughly investigated the trilayered 6-state clock model by Monte Carlo simulation. The results are reported with diagrams in this manuscript. This manuscript is organized as follows: The model and the simulational scheme are discussed in Section-2. The numerical results are reported with diagrams in Section-3 and the paper ends with concluding remarks in Section-4.
@@ 2 Model and numerical details
We consider a classical trilayered 6-state clock model defined on a simple cubic lattice, where each of the three monoatomic layers—denoted $l_{1}$, $l_{2}$, $l_{3}$—consists exclusively of either type-A or type-B atoms. The magnetic interactions and anisotropic contributions to the energy of the system are captured by the following Hamiltonian:
@@
Here, $\vec{S_{i}}=(S_{i}^{x},S_{i}^{y})$ represents the classical two dimensional spin vector at lattice site $i$, which can point in one of six discrete directions in the $xy$-plane, consistent with the 6-state clock model. These discrete orientations are defined by:
@@
The first term in Eq-(1) accounts for the intralayer ferromagnetic interactions between nearest neighbour spins $\langle ij\rangle$ within the same atomic plane, with a positive coupling exchange $J_{ferro}^{intra}>0$. The second term represents interlayer antiferromagnetic coupling between the adjacent layers, involving nearest neighbour sites $\langle ij\rangle$ across the layers, governed by a coupling constant $J_{antiferro}^{inter}<0$. The third term introduces an anisotropy D, which destroys the discrete rotational symmetry in the $xy$-plane and favours the spin-alignment in specific direction. Positive values of $D$ prefer specific spin orientations, thereby reinforcing the discrete six-fold symmetry and enhancing the stability of the selected clock states. We impose periodic boundary conditions (PBC) in the x and y directions within each layer and along the stacking direction z; we adopt free boundary conditions to capture the finite thickness of the trilayered structure.
To study the compensation behaviour in the trilayered 6-state clock model, we employ Monte Carlo simulation technique based on Metropolis algorithm with random single-spin updates. The system consists of three stacked square lattices, each consisting of $L\times L$ sites, resulting in a total of $3L^{2}$ spins. For majority of our simulations, we consider a system size of $L=128$, corresponding to $128\times 128$ spins per layer. In order to examine finite-size effects, additional simulations are performed for various lattice sizes. Initially, spin configurations are generated at high temperature with randomly assigned orientations chosen from the six allowed states in the clock model. Then the system is allowed to evolve using Metropolis updating scheme: during each Monte Carlo steps (MCS), $L^{2}$ spin-flip attempts are made, where a trial spin selected at random and updated according to the Metropolis acceptance probability[27]: $P_{f}=Min[\exp{-\frac{\delta E}{k_{B}T}},1]$. $\Delta E$ is the change in energy associated with the proposed spin flip, and $k_{B}(=1)$ is the Boltzmann constant. The simulation spans a duration $2\times 10^{5}$ MCSS within an initial transient segment of $10^{5}$ MCSS excluded to achieve the assumed ergodic limit. To characterize the thermodynamic behaviour of the 6-state clock model, we calculate several key observables that capture the collective magnetic ordering of the system. The sublattice magnetisation provides insight into the local symmetry breaking and is defined for each sublattice $k\in{1,2,3}$ as the average spin components in the x and y directions:
@@
where $S^{x}(i,j,k)$ and $S^{y}(i,j,k)$ denote the Cartesian components of the spin vector on the site $(i,j)$ belonging to the $k$-th sublattice and $L^{2}$ is the total number of lattice site in each sublattice. The total magnetization of the system is obtained by summing the contributions from all sublattices:
@@ 3.1 Isotropic regime:$D=0$
First, we discuss the thermodynamic phase transition and compensation phenomenon of a trilayered 6-state clock at zero anisotropy($D=0$). All six clock states are energetically favourable in this regime, as there is no directional bias. The intralayer ferromagnetic coupling is taken for A-type atoms as $J_{ferro}^{AA}=0.2$ and for B-type atoms as $J_{ferro}^{BB}=1.0$. The interlayer antiferromagnetic coupling strength is taken as $J_{antiferro}^{AB}=-0.1$. Initially, the system is at high-temperature disordered paramagnetic state, characterised by random spin orientations and zero magnetisation. The system is slowly cooled down with small temperature step $dT=0.05$. As the temperature decreases, magnetic ordering sets in. Figure-1(a) shows the growth of total magnetization $M(T)$, which increases continuously from zero to a non-zero finite value at a well-defined critical temperature $T_{c}$. The transition temperature $T_{c}$ is precisely identified from the position of the peak of the magnetic susceptibility $\chi(T)$ as shown in Figure-1(b).
@@
Interestingly, below $T_{c}$, the total magnetisation again vanishes at a lower temperature $T_{comp}$, referred to as compensation temperature. This occurs even though the sublattices (individual layers) remain partially ordered. Compensation is a peculiar balancing situation where the total magnetization of all layers vanishes, despite the nonzero sublattice ordering. The susceptibility curve $\chi(T)$ exhibits a secondary peak at this temperature, further supporting the presence of a sharp magnetic cross-over. The corresponding specific heat $C(T)$ shown in Figure-1(c), also reflects these thermodynamic features with distinct peaks at both $T_{c}$ and $T_{comp}$. The compensation phenomenon originates from the competing magnetic interations: each layer offers ferromagnetic intra-layer coupling, favouring uniform spin alignment within the layer, where adjacent layers are coupled via anti-ferromagnetic interlayer interactions leading to repel layer magnetisation.
@@
To acquire a more nuanced perspective to the magnetic ordering in each layer, we analyse the sublattice magnetisation of the individual layers across the temperature. Figure-2 shows the sublattice magnetisation components $m_{x}$ and $m_{y}$ as functions of temperature for three distinct layers. These profiles describe the growth of magnetic ordering in each layer. At $D=0$, where no easy-axis anisotropy is present, the system does not energetically favour in any specific direction in the $xy$ plane. Consequently, both $m_{x}$ and $m_{y}$ develop finite values below the transition temperature, indicating spontaneous symmetry breaking. As shown in Figure-2(a) or (b) all three layers exhibit ferromagnetic ordering below the transition temperature $T_{c}$ with each layer developing finite magnetisation. However, due to the antiferromagnetic interlayer coupling, the magnetisations of adjacent layers are oriented in opposite directions. In particular, the middle layer tends to be antiparallel to the outer layers, leading to cancellation of total magnetisation at $T_{comp}$.
@@
To visualise the spin ordering, we also examine the lattice morphology at representative temperatures. Spin configurations snapshots displayed in Figure-3 reveal a transition from a disordered paramagnetic state at high temperature to a well-aligned domain at low temperature for different layers. At very high temperature $T>>T_{c}$ spin are randomly oriented in all three layers. Six colours denote the six specific clock states. For the isotropic regime $(D=0)$, at high temperature all six directions are equally preferable. As temperature decreases,the spins at each layer progressively align, leading to the nucleation and growth of the magnetic domain, which marks the onset of long-range order and results a non-zero total magnetization. Near $T_{comp}$, distinct domain patterns emerge with opposing magnetisation contributions, reflecting the layer-wise compensation behaviour. To quantatively validate the spin morphology, we analysed the angular distribution of spins across the three layers at different temperature. The normalized spin population (fraction of total spins) is plotted as a function of six discrete clock states ( corresponding spin angle 0, $\pi/3$, $2\pi/3$, $3\pi/3$, $4\pi/3$, $5\pi/3$) for each layer. At high temperature, the spin population is nearly uniform across all angles in all three layers, reflecting an isotropic regime where all spin states are equally probable. As temperature decreases towards the critical value, certain spin angles become more populated, indicating the onset of preferred aligment. Near compensation temperature, layer 1 and 3 exihibit a dominant population at $\theta=\pi/3$(n=2), whereas layer 2 prefers $\theta=5\pi/3$. Importantly, the other spin angles are populated such that net magnetization of the whole system becomes zero, while each layer individually retains a finite magnetisation. Upon further cooling to low temperature, layers 1 and 3 show nearly complete alignment at $\theta=\pi/3$ while layer 2 remains oriented in $5\pi/3$. The visual and quantitative results together confirm the layered character of magnetic ordering and the competing interactions that drive the compensation phenomenon.

### v2

@@ Abstract
The equilibrium behaviours of a trilayered 6-state clock model has been investigated by Monte Carlo simulation. The intralayer interaction is considered ferromagnetic, whereas the interlayer interaction is antiferromagnetic. Periodic boundary conditions(PBC) are applied in XY plane and open boundary condition (OBC) is applied along Z-direction. The thermodynamic behaviours of sublattice magnetisation, total magnetisation, magnetic susceptibility and the specific heat are studied as functions of the temperature. The interesting compensation phenomenon, i.e., the vanishing of the total magnetisation even with nonvanishing sublattice magnetisation, has been observed. The compensation temperature and the critical temperature are found to depend on the strength of single-site anisotropy. The comprehensive phase diagram is shown in the temperature-anisotropy plane. The finite size analysis has been done through the temperature dependence of fourth order Binder cumulant, and the scaling exponents are estimated.
@@ 1 Introduction
The ferromagnetic models showing the unusual ordered phase even in absence of conventional long-range ferromagnetic order is a remarkable discovery[1, 2]. The continuous symmetric (SO(2)) planar ferromagnetic model shows this behaviour in two dimensions with vortex-antivortex special kind of ordered state[3, 4, 5]. The unconventional metastable behaviour and phase transition in this kind of model has drawn the attention of the researchers in last four decades. The celebrated BKT transition is now under the spotlight of modern research in condensed matter physics and statistical mechanics.
@@
Is this kind of unusual phase the property of continuous symmetry only? This question has triggered to investigate the discrete symetric q-state clock model, where the continuous symetric XY model can be restored in the limit $q\to\infty$. The properties of q-state clock model has been investigated[6] for q=4, 5 and 6 in two dimensions by Monte Carlo renormalization group analysis to study the phase transitions. The Kosterlitz-Thouless type intermediate phase has been found[7] in two dimensional 6-state clock model by Monte Carlo simulations. The KT type phase transition has been found to occur in the range between ${{k_{B}T}\over{J}}=0.68\pm 0.02$ and ${{k_{B}T}\over{J}}=0.92\pm 0.01$. However, it is confirmed later that there exists a Kosterlitz-Thouless (KT) like phase in the ferromagnetic 6-clock model on the square lattice through studies of the interfacial free energy estimated by Monte Carlo simulations and the authors estimated[8] that lower and upper transition temperatures are 0.75 and 0.90 respectively. Interestingly, in a recent study[9], the machine-learning detection of BKT transition was found in q-state clock model with the neural network based estimates of the two critical points as ${{k_{B}T}\over{J}}=0.410$ and ${{k_{B}T}\over{J}}=0.921$.
@@
The classical 2D clock model is known to have a critical phase with Berezinskii–Kosterlitz–Thouless (BKT) transitions. These transitions have logarithmic corrections which make numerical analysis difficult. Hence, the classical 2D six-state clock model has been studied [10] by mapping it to the quantum self-dual 1D six-state clock model and confirmed that the self-dual point has a precise numerical agreement with the analytical result. It has also been argued that the degeneracy of the excited states at the self-dual point from the effective field theoretical point of view.
@@
The thermodynamic behaviours of the clock model has been studied in different kinds of lattices to investigate the effects of geometrical connections and interactions. The BKT transition has been observed[11] in 6-state clock model on the rewired squared lattice studied by Monte Carlo simulation with Wang-Landau algorithm and estimated the lower and upper critical temperatures. For pure case, i.e., the lattice with average number of bonds equal to 2, ${k_{B}T1}\over{J}$} and ${k_{B}T2}\over{J}$}, are roughly estimated around 0.55 and 1.15 respectively (these results may be compared to those reported in the references [7] [8] and [9]). The BKT transition has been studied[12] recently, in 6-state clock model on different kinds (triangular, kagome, honeycomb and diced) of two-dimensional lattices. The two-dimensional q-state clock model on random lattices has been investigated clock-random via numerical simulation. Interestingly, for q=4 a single phase transition has been reported. Whereas, for q=5 and q=6 multiple (double) phase transitions are observed from the two peaks of the specific heat. The lower and upper critical temperatures are found to depend significantly on the structure of the lattice. The anisotropic (with anisotropy parameter $\alpha$) 6-state clock model has been investigated[14] by transfer matrix renormalization group method to explore the phase transition in two dimensional square lattice. Three distinct phase are identified and a comprehensive phase diagram is drawn. They observed a single discontinuous phase transition and two BKT phase transitions meeting in the tricritical point. The tricritical point is estimated at critical anisotropy $\alpha_{c}=0.21405$. The antiferromagnetic 5-state clock-model has been studied on the triangular lattice recently[15] with second nearest neighbour interaction and the phase diagram is drawn.
@@
The thermodynamic behaviours of clock model in layered structured lattices are not yet reported in the well developed literature of this specific model. However, the behaviours of layered Ising and Blume-Capel models are well studied. Let us briefly mention a few selected studies on layered Ising and Blume-Capel models. The main interesting phenomenon in the thermodynamic behaviours of layered Ising and Blume-Capel models is the compensation transition. The compensation temperatures and magnetic susceptibility of a mixed ferro-ferrimagnetic ternary alloy have been studied[16] by Monte Carlo simulations and observed outstanding behaviour of magnetic susceptibility near the compensation and critical temperature. The compensation temperature has been found[17] to depend of the anisotropy in a mixed spin (spin-${1}\over{2}$} and spin-$1$) Heisenberg ferrimagnet on a simple cubic lattice, studied by Oguchi approximation. The anisotopy dependent spin compensation temperature has also been noticed[18] in mixed spin (spin-${3}\over{2}$} and spin-${5}\over{2}$}) ferrimagnetic Ising model studied by mean field approximation. In this case, the possibility of multiple compensation temperature was indicated. The impurity- concentration dependent compensation[19] and the anisotropy dependent compensation[20] was also noticed in the trilayered Ising and Blume-Capel models, respectively. The meanfield and effective field theory has been employed [21] to study the interaction dependent compensation in spin-${1}\over{2}$} Ising trilayered ferrimagnetic system. The phase diagram (compensation and no-compensation) was drawn in parameter space. The antiferromagnetic systems showed the compensation behaviours in Ising mixed spin (spin-${3}\over{2}$} and spin-${5}\over{2}$}) core/shell nanowire[22] and in mixed spin (spin-${7}\over{2}$} and spin-1{1}) antiferromagnetic ovalene nanowire[23]. The compensation wad found[24] Ising mixed spin (spin-${3}\over{2}$} and spin-${1}\over{2}$}) graphene bilayered ferrimagnet. The effects of exchange interactions, crystal field and external magnetic field on the compensation temperature were studied [24] by Monte Carlo simulation. The evidence of engineering the compensation in ferrimagnetic iron garnet by tuning the ratio of in-plane and out of plane magnetisations can be found in [25]. An interesting spin compensation and first-order reentrant behaviour has been observed[26] in mixed spin (spin-1 and spin-2) Ising ferrimagnetic system studied by Monte Carlo simulation.
@@
Can the discrete symmetric planar models (6-state clock model) show the compensation behaviour? Being motivated by this question, we have thoroughly investigated the trilayered 6-state clock model by Monte Carlo simulation. The results are reported with diagrams in this manuscript. This manuscript is organised as follows: The model and the simulation scheme are discussed in Section-2. The numerical results are reported with diagrams in Section-3 and the paper ends with concluding remarks in Section-4.
@@ 2 Model and numerical details
We consider a classical trilayered 6-state clock model defined on a simple cubic lattice, where each of the three monoatomic layers consists exclusively of either type-A or type-B atoms. The magnetic interactions and anisotropic contributions to the energy of the system are captured by the following Hamiltonian:
@@
Here, $\vec{S_{i}}=(S_{i}^{x},S_{i}^{y})$ represents the classical two-dimensional spin vector at lattice site $i$, which can point in one of six discrete directions in the $xy$-plane, consistent with the 6-state clock model. These discrete orientations are defined by:
@@
The first term in Equation-1 accounts for the intralayer ferromagnetic interactions between nearest neighbour spins $\langle ij\rangle$ within the same atomic plane, with a positive coupling exchange $J_{ferro}^{intra}>0$. The second term represents interlayer antiferromagnetic coupling between the adjacent layers, involving nearest neighbour sites $\langle ij\rangle$ across the layers, governed by a coupling constant $J_{antiferro}^{inter}<0$. The third term introduces an anisotropy D, which destroys the discrete rotational symmetry in the $xy$-plane and favours the spin-alignment in specific direction. Positive values of $D$ prefer specific spin orientations, thereby reinforcing the discrete six-fold symmetry and enhancing the stability of the selected clock states. We impose periodic boundary conditions (PBC) in the x and y directions within each layer and along the stacking direction z; we adopt free boundary conditions to capture the finite thickness of the trilayered structure.
To study the compensation behaviour in the trilayered 6-state clock model, we employ the Monte Carlo simulation technique based on the Metropolis algorithm with random single-spin updates. The system consists of three stacked square lattices, each consisting of $L\times L$ sites, resulting in a total of $3L^{2}$ spins. For the majority of our simulations, we consider a system size of $L=128$, corresponding to $128\times 128$ spins per layer. In order to examine finite-size effects, additional simulations are performed for various lattice sizes. Initially, spin configurations are generated at high temperature with randomly assigned orientations chosen from the six allowed states in the clock model. Then the system is allowed to evolve using Metropolis updating scheme: during each Monte Carlo steps (MCS), $L^{2}$ spin-flip attempts are made, where a trial spin selected at random and updated according to the Metropolis acceptance probability[27]: $P_{f}=Min[\exp{-\frac{\delta E}{k_{B}T}},1]$. $\Delta E$ is the change in energy associated with the proposed spin flip, and $k_{B}(=1)$ is the Boltzmann constant. The simulation spans a duration $3\times 10^{5}$ MCSS within an initial transient segment of $1.5\times 10^{5}$ MCSS excluded to achieve the assumed ergodic limit. To characterize the thermodynamic behaviour of the 6-state clock model, we calculate several key observables that capture the collective magnetic ordering of the system. The sublattice magnetisation provides insight into the local symmetry breaking and is defined for each sublattice $k\in{1,2,3}$ as the average spin components in the x and y directions:
@@
where $S^{x}(i,j,k)$ and $S^{y}(i,j,k)$ denote the Cartesian components of the spin vector on the site $(i,j)$ belonging to the $k$-th sublattice and $L^{2}$ is the total number of lattice sites in each sublattice. The total magnetisation of the system is obtained by summing the contributions from all sublattices:
@@ 3.1 Isotropic regime:$D=0$
First, we discuss the thermodynamic phase transition and compensation phenomenon of a trilayered 6-state clock model at zero anisotropy($D=0$). All six clock states are energetically favourable in this regime, as there is no directional bias. The intralayer ferromagnetic coupling is taken for A-type atoms as $J_{ferro}^{AA}=0.2$ and for B-type atoms as $J_{ferro}^{BB}=1.0$. The interlayer antiferromagnetic coupling strength is taken as $J_{antiferro}^{AB}=-0.1$. Initially, the system is at high-temperature disordered paramagnetic state, characterised by random spin orientations and zero magnetisation. The system is slowly cooled down with small temperature step $dT=0.05$. As the temperature decreases, magnetic ordering sets in. Figure-1(a) shows the growth of total magnetisation $M(T)$, which increases continuously from zero to a non-zero finite value at a well-defined critical temperature $T_{c}$. The transition temperature $T_{c}$ is precisely identified from the position of the peak of the magnetic susceptibility $\chi(T)$ as shown in Figure-1(b).
@@
Interestingly, below $T_{c}$, the total magnetisation again vanishes at a lower temperature $T_{comp}$, referred to as compensation temperature. This occurs even though the sublattices (individual layers) remain partially ordered. Compensation is a peculiar balancing situation where the total magnetisation of all layers vanishes, despite the nonzero sublattice ordering. The susceptibility curve $\chi(T)$ exhibits a secondary peak at this temperature, further supporting the presence of a sharp magnetic crossover. The corresponding specific heat $C(T)$ shown in Figure-1(c), also reflects these thermodynamic features with distinct peaks at both $T_{c}$ and $T_{comp}$. The compensation phenomenon originates from the competing magnetic interactions: each layer offers ferromagnetic intra-layer coupling, favouring uniform spin alignment within the layer, where adjacent layers are coupled via anti-ferromagnetic interlayer interactions leading to repel layer magnetisation.
@@
To acquire a more nuanced perspective on the magnetic ordering in each layer, we analyse the sublattice magnetisation of the individual layers across the temperature. Figure-2 shows the sublattice magnetisation components $m_{x}$ and $m_{y}$ as functions of temperature for three distinct layers. These profiles describe the growth of magnetic ordering in each layer. At $D=0$, where no easy-axis anisotropy is present, the system does not energetically favour in any specific direction in the $xy$ plane. Consequently, both $m_{x}$ and $m_{y}$ develop finite values below the transition temperature, indicating spontaneous symmetry breaking. As shown in Figure-2(a) or (b) all three layers exhibit ferromagnetic ordering below the transition temperature $T_{c}$ with each layer developing finite magnetisation. However, due to the antiferromagnetic interlayer coupling, the magnetisations of adjacent layers are oriented in opposite directions. In particular, the middle layer tends to be antiparallel to the outer layers, leading to cancellation of total magnetisation at $T_{comp}$.
@@
To visualise the spin ordering, we also examine the lattice morphology at representative temperatures. Spin configurations snapshots displayed in Figure-3 reveal a transition from a disordered paramagnetic state at high temperature to a well-aligned domain at low temperature for different layers. At very high temperature $T>>T_{c}$ spin are randomly oriented in all three layers. Six colours denote the six specific clock states. For the isotropic regime $(D=0)$, at high temperature, all six directions are equally preferable. As temperature decreases, the spins at each layer progressively align, leading to the nucleation and growth of the magnetic domain, which marks the onset of long-range order and results a non-zero total magnetisation. Near $T_{comp}$, distinct domain patterns emerge with opposing magnetisation contributions, reflecting the layer-wise compensation behaviour. To quantitatively validate the spin morphology, we analysed the angular distribution of spins across the three layers at different temperatures. The normalised spin population (fraction of total spins) is plotted as a function of six discrete clock states ( corresponding spin angle 0, $\pi/3$, $2\pi/3$, $3\pi/3$, $4\pi/3$, $5\pi/3$) for each layer. At high temperature, the spin population is nearly uniform across all angles in all three layers, reflecting an isotropic regime where all spin states are equally probable. As temperature decreases towards the critical value, certain spin angles become more populated, indicating the onset of preferred alignment. Near compensation temperature, layer 1 and 3 exhibit a dominant population at $\theta=\pi/3(n=2)$, whereas layer 2 prefers $\theta=5\pi/3(n=6)$. Importantly, the other spin angles are populated such that net magnetisation of the whole system becomes zero, while each layer individually retains a finite magnetisation. Upon further cooling to low temperature, layers 1 and 3 show nearly complete alignment at $\theta=\pi/3$ while layer 2 remains oriented in $5\pi/3$. The visual and quantitative results together confirm the layered character of magnetic ordering and the competing interactions that drive the compensation phenomenon.